# Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model

### v1

@@ 4.2 Implementation
We employ GPT-4 as the teacher LLM to construct our training dataset and utilize FlanT5 models—Small (60M), Base (250M), and Large (760M) (Chung et al., 2022)—as student SLMs. We manually create 8 demonstrations to guide GPT-4 in generating 4 reasoning paths for each dataset (KPDD-CoT and KPDD-PoT). Fine-tuning of all student SLMs is conducted using the Huggingface library (Wolf et al., 2020) on an NVIDIA 3090 GPU with 24 GB RAM. The learning rate for fine-tuning is set to 5e-4, with a total of 10 fine-tuning epochs.
4.3 Baselines
Proprietary Large Language Models We present CoT prompting results from an array of SoTA LLMs, such as OpenAI’s GPT-4, ChatGPT (gpt-3.5-turbo), Google’s PaLM-2, and Anthropic’s Claude-2.
Open-Source Large Language Models We present mathematical reasoning performance of Llama-2-7B, CodeLLaMA-7B, and their fine-tuned versions, such as Platpus-2, WizardMath, TORA.
Fine-tuned Small Language Models We present some works that try to fine-tune SLMs under 1B, such as Ho et al. Ho et al. (2023) fine-tune GPT-3-ada, Fu et al. Fu et al. (2023) fine-tune FlanT5, and Shridhar et al. Shridhar et al. (2023) fine-tune GPT-2.
@@ 4.3 Baselines
Table 2: Overall test set performance. We use KPDD to fine-tune SLMs, and evaluate them on four mathematical reasoning datasets, i.e., GSM8K, ASDiv, SVAMP, and MultiArith. The experiment results show that KPDD-CoT can effectively improve SLMs’ reasoning performance, and KPDD-PoT makes SLMs achieve SOTA reasoning performance.
@@ 4.4 Main Results
    KPDD-CoT Enhances Mathematical Reasoning: KPDD-CoT significantly improves the mathematical reasoning capabilities of SLMs, with absolute improvements ranging from 5.01% to 15.51% across tasks. Traditional baselines typically rely on CoTD, which involves generating numerous steps and performing extensive calculations. However, CoTD often encounters semantic misunderstanding errors that hinder the improvement of SLMs’ mathematical reasoning abilities. In contrast, KPDD-CoT employs extra SLMs to extract key points (including the core question and problem-solving information) of the question and uses these key points to guide the SLMs’ reasoning. This approach significantly reduces the semantic misunderstanding errors of CoTD, making KPDD-CoT better suited for improving the mathematical reasoning ability of SLMs.
@@
    KPDD-PoT Outperforms State-of-the-Art: KPDD-PoT surpasses previous state-of-the-art fine-tuned SLMs at all scales, with absolute improvements between 32.18% and 54.63% across tasks. Furthermore, KPDD-PoT’s accuracy is higher than that of KPDD-CoT, highlighting the advantage of rationales in PoT format in enhancing SLMs’ reasoning capabilities. Our analysis finds that the mathematical reasoning performance of CoTD is limited not only by semantic misunderstanding errors but also by calculation errors. PoTD converts rationales from CoT format into PoT format, formulating the reasoning process into a Python program and sending it to an extra Python interpreter to generate the final answer. This method transfers numerical computation from SLMs to a Python interpreter, avoiding calculation errors. Additionally, by extracting key points of the question, KPDD-PoT implicitly enhances the SLMs’ understanding of the question, thereby improving their overall mathematical reasoning capabilities.
@@
    Importance of Model Size: The efficacy of mathematical reasoning distillation in SLMs is highly dependent on model size; larger models assimilate more reasoning knowledge, leading to superior performance. For instance, under KPDD-PoT, FlanT5-Small achieves 20.77% accuracy on GSM8K, FlanT5-Base reaches 34.57%, and FlanT5-Large attains 46.32%.
  4. 4.
@@ 4.7 Diverse Reasoning Paths Improve SLMs’ Reasoning Performance
In this subsection, we fine-tune CodeT5-Base on our reasoning datasets, which are differentiated by the number of reasoning paths they contain, to analyze the effect of reasoning path multiplicity on the reasoning performance of SLMs. This examination aims to discern how the quantity of reasoning paths in training data influence the model’s ability to perform reasoning tasks.
@@
Figure 5 presents the results of our experiments, which demonstrate that a variety of reasoning paths can bolster the reasoning performance of SLMs. For instance, CodeT5-Base, when trained on an KPDD-PoT dataset featuring four reasoning paths, attains a 34.57% accuracy on the GSM8K test dataset and a 52.29% accuracy on ASDiv. In contrast, CodeT5-Base trained on an KPDD-PoT dataset with only one reasoning path achieves 49.33% accuracy on GSM8K test dataset and 46.1% accuracy on ASDiv. This suggests that the inclusion of multiple reasoning paths in training data can significantly enhance the model’s performance, particularly in tasks requiring explanation generation.

### v2

@@ 4.2 Implementation
In our paper, we use GPT-4 as our teacher LLM to generate reasoning processes for questions. Specifically, we manually construct a demonstration set containing eight demonstrations. We then use this demonstration set to prompt GPT-4 to generate four reasoning paths for each question. Each reasoning path is subsequently filtered, resulting in the creation of a KPDD dataset. Next, we use FlanT5 models—Small (60M), Base (250M), and Large (760M) (Chung et al., 2022)—as our student LMs. By fine-tuning the FlanT5 models with the KPDD dataset, we aim to enhance their mathematical reasoning abilities. During the fine-tuning process, we set the learning rate to 5e-4, the batch size to 32, and the total number of training epochs to 10.
@@
Table 2: Overall test set performance.
@@ 4.3 Main Results
    KPDD-CoT Enhances Mathematical Reasoning: When FlanT5-small is used as the SLM, KPDD-CoT achieves an average accuracy improvement of 5.01% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-CoT yields an average accuracy improvement of 11.71%. For FlanT5-large, KPDD-CoT results in an average accuracy improvement of 15.51%. The experimental result demonstrates that KPDD-CoT can significantly enhance the mathematical reasoning performance of SLMs. We attribute this experimental result to the reason that baselines often encounters semantic misunderstanding errors that hinder the improvement of SLMs’ mathematical reasoning abilities. In contrast, KPDD-CoT employs extra SLMs to extract key points (including the core question and problem-solving information) of the question and uses these key points to guide the SLMs’ reasoning. This approach significantly reduces the semantic misunderstanding errors of CoTD, making KPDD-CoT better suited for improving the mathematical reasoning ability of SLMs.
@@
    KPDD-PoT Outperforms State-of-the-Art: When FlanT5-small is used as the SLM, KPDD-PoT achieves an average accuracy improvement of 32.18% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-PoT yields an average accuracy improvement of 48.25%. For FlanT5-large, KPDD-PoT results in an average accuracy improvement of 54.63%. The experimental result shows that KPDD-PoT make SLMs achieve state-of-the-art mathematical reasoning accuracy. Furthermore, KPDD-PoT’s accuracy is higher than that of KPDD-CoT, highlighting the advantage of rationales in PoT format in enhancing SLMs’ reasoning capabilities. Our analysis finds that the mathematical reasoning performance of CoTD is limited not only by semantic misunderstanding errors but also by calculation errors. PoTD converts rationales from CoT format into PoT format, formulating the reasoning process into a Python program and sending it to an extra Python interpreter to generate the final answer. This method transfers numerical computation from SLMs to a Python interpreter, avoiding calculation errors. Additionally, by extracting key points of the question, KPDD-PoT implicitly enhances the SLMs’ understanding of the question, thereby improving their overall mathematical reasoning capabilities.
@@ 4.6 Diverse Reasoning Paths Improve SLMs’ Reasoning Performance
In this subsection, we primarily explore the impact of the diversity of reasoning paths on the mathematical reasoning performance of SLMs. Specifically, we fine-tune FlanT5-base using KPDD datasets with varying numbers of reasoning paths and then evaluate the fine-tuned SLMs on several mathematical reasoning datasets. By analyzing their mathematical reasoning performance, we assess how the diversity of reasoning paths affects the SLMs’ capabilities in mathematical reasoning.
@@
Figure 5 reveals the relevant experimental results. As the number of reasoning paths in the distillation dataset increases, the SLM tends to achieve better performance. For example, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing a single reasoning path, it achieved an accuracy of 9.17% on the GSM8K test dataset, 9.92% on ASDiv, 9.03% on SVAMP, and 11.83% on MultiArith. However, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing four reasoning paths, it achieved accuracies of 14.63% on the GSM8K test dataset, 14.93% on ASDiv, 13.8% on SVAMP, and 21.5% on MultiArith. Similarly, when fine-tuning SLMs with KPDD-PoT datasets containing different reasoning paths, the experimental results were consistent with those of KPDD-CoT. These results indicate that a mathematical reasoning dataset with diverse reasoning paths can effectively enhance the mathematical reasoning performance of SLMs.